# The Instrumental Background of EP/FXT

### v1

@@ 6 Orbit Modulation
The entire instrumental background rate distribution on the orbit is plotted in the bottom panel of Figure 6, where all the SCD in thin filters and the FWC data are mapped onto the geographical coordinate grids. This panel displays the distribution with enhanced clarity. Apart from the edge of the SAA region, the rate increases towards the north geomagnetic pole direction, i.e. the region near (30N, 90W), and around the mid-latitudes in the south. This pattern is directly correlated with the geomagnetic field, the cut-off rigidity (Smart & Shea, 2009; Gerontidou et al., 2021) and the LEO proton and electron flux distributions (Koshiishi, 2014).
@@
To examine the spectral variations at varying intensities, we classify the instrumental background events of the FXTA FWC observation into three groups based on the count rate value: rate¡0.4 counts/s, $\in$[0.4,0.6] counts/s and rate¿0.6 counts/s. The spectra corresponding to these three groups are depicted in the top panel of Figure 7. The normalization was adjusted to facilitate a comparison among the spectral shapes in the lower panel. It could be concluded that the instrumental background spectral shape in the 0.5–10 keV energy band remains consistent across different count rates, i.e. at varying geomagnetic fields. Similar features have been observed in the silicon semiconductor detector of LE telescope aboard the Insight-HXMT mission (Liao et al., 2020). The analysed spectra in Figure 5 also demonstrate that the instrumental background retains similar morphology and flux levels within one year period, showing no significant variation.
@@ 8 Long-term Variation
Figure 10 also indicates that the daily IMG instrumental background rate during the initial $\sim$ 200 days appears higher than in subsequent epochs. This decreases might relate with the change of the space environments. The year of 2025 is around solar maximum. Theoretically, the local interstellar Galactic cosmic rays at lower energies, usually at $<10$ GeV, entering the solar system will become less due to the higher potential during the solar maximum years (Zhang et al., 2022, and reference therein). Therefore, the instrumental background rate would exhibit an anti-correlation with the solar activity from a long-term perspective, as viewed by XMM-Newton/EPIC-pn (Bulbul et al., 2020) and Chandra/ACIS (Suzuki et al., 2021; Grant et al., 2022) based on data accumulated over decade years. As also plotted in Figure 10 is the monthly averaged 10.7 cm radio flux, which serves as a proxy for solar activity. Currently, the particle-induced data collected over one year on FXT have not demonstrated this significant correlation.
As the reduction in orbit height also leads to the decrease in instrumental background rate, due to the shielding effect of the Earth magnetic field against low energy primary cosmic rays, the daily averaged EP altitude is additionally dotted on Figure 10 for comparison. So far the altitude has degraded approximately 20 km since launch. And the secular instrumental background rate reduced with time at a speed of $(5.44\pm 0.14)\times 10^{-5}$ counts/s per day. The orbit decay could potentially account for the secular reduction. However, further long-term data accumulation is required to analyze the impact of the solar activity on FXT instrumental background.
@@ 9 Summary and Conclusion
Distinct fluorescence lines from Al, Fe, Ni, and Cu are clearly observed in the instrumental background spectra. These lines could be employed for the in-orbit energy calibration. The spectral shape at 0.5–10 keV remains consistent at different rate regimes. Furthermore, the obtained FSA rate displays a good linear correlation with the IMG, thanks to the innovative FF readout design of FXT. Based on these findings, the IMG instrumental background modelling was established, and validated by the FWC and the Lockman Hole observations. This modelling could facilitate the analysis of extended source observations.
Acknowledgements.

### v2

@@ 6 Orbital Modulation
Figure 6: The instrumental background variation with time and geographical location. (a) The light curve in 0.5–10 keV of one FXTA FWC observation. (b) The corresponding geography distribution of this light curve. (c) The geography distribution of all SCD in thin filter and FWC instrumental background observations. The color bars in Panels (b) and (c) are labeled in units of counts/s, and the SAA region utilized for EP/FXT is delineated by the pink dashed line. Please refer to the text for more detailed descriptions.
@@
The entire instrumental background rate distribution on the orbit is plotted in the bottom panel of Figure 6, where all the SCD in thin filters and the FWC data are mapped onto the geographical coordinate grids. This panel displays the distribution with enhanced clarity. Apart from the edge of the SAA region, the rate increases towards the north geomagnetic pole direction, i.e. the region near (30N, 90W), and around the mid-latitudes in the south. This pattern is directly correlated with the geomagnetic field, the cut-off rigidity (D. F. Smart & M. A. Shea, 2009; M. Gerontidou et al., 2021) and the LEO proton and electron flux distributions (H. Koshiishi, 2014).
@@
To examine the spectral variations at varying intensities, we classify the instrumental background events of the FXTA FWC observation into three groups based on the count rate value: rate<0.4 counts/s, $\in$[0.4,0.6] counts/s and rate>0.6 counts/s. The spectra corresponding to these three groups are depicted in the top panel of Figure 7. The normalization was adjusted to facilitate a comparison among the spectral shapes in the lower panel. It could be concluded that the instrumental background spectral shape in the 0.5–10 keV energy band remains consistent across different count rates, i.e. at varying geomagnetic fields. Similar features have been observed in the silicon semiconductor detector of LE telescope aboard the Insight-HXMT mission (J.-Y. Liao et al., 2020). The analysed spectra in Figure 5 also demonstrate that the instrumental background retains similar morphology and flux levels within one year period, showing no significant variation.
@@ 8 Long-term Variation
Figure 10 also indicates that the daily IMG instrumental background rate during the initial $\sim$ 200 days appears higher than in subsequent epochs. This decrease might be attributed to the fact that in 2024 the Sun was evolving towards its maximum of activity, expected around mid 2025, then the shielding of cosmic rays of the solar potential below a few GeV since launch became more and more efficient. Therefore, the instrumental background rate would exhibit an anti-correlation with the solar activity from a long-term perspective, as viewed by XMM-Newton/EPIC-pn (E. Bulbul et al., 2020) and Chandra/ACIS (H. Suzuki et al., 2021; C. E. Grant et al., 2022) based on data accumulated over decade years. As also plotted in Figure 10 is the monthly averaged 10.7 cm radio flux, which serves as a proxy for solar activity. Currently, the particle-induced data collected over one year on FXT have not significantly demonstrated this correlation due to the fluctuation of the solar activity. Further long-term data accumulation is required to analyze the impact of the solar activity on FXT instrumental background.
As the reduction in orbit height also leads to the decrease in instrumental background rate, due to the shielding effect of the Earth magnetic field against low energy primary cosmic rays, the daily averaged EP altitude is additionally dotted on Figure 10 for comparison. Till the end of May 2025, the altitude has degraded approximately 20 km since launch. And the secular instrumental background rate reduced with time at a speed of $(5.44\pm 0.14)\times 10^{-5}$ counts/s per day. The orbit decay could partially account for the secular reduction.
@@ 9 Summary and Conclusion
Distinct fluorescence lines from Al, Fe, Ni, and Cu are clearly observed in the instrumental background spectra. These lines could be employed for the in-orbit energy calibration. The spectral shape in the 0.5–10 keV band remains consistent at different rate regimes. Furthermore, the obtained FSA rate displays a good linear correlation with the IMG, thanks to the innovative FF readout design of FXT. Based on these findings, the IMG instrumental background modelling was established, and validated by the FWC and the Lockman Hole observations. This modelling could facilitate the analysis of extended source observations such as those of galaxy clusters.